# Think Before You Diffuse: LLMs-Guided Physics-Aware Video Generation

### v1

@@ Abstract
Figure 1: DiffPhy enables physically grounded and semantically aligned video generation across diverse real-world scenarios, including gravity-driven motion, fluid interactions, forceful impacts, and object manipulation. It outperforms state-of-the-art video diffusion model Wan 2.1–14B [50] in both visual plausibility and physical coherence.
@@ 1 Introduction
Recent advances in video diffusion models have made it possible to generate high-fidelity videos with rich details from text prompts [15, 50, 63, 10, 27, 2, 40, 35, 52, 13, 42, 45, 26]. Such generation capability has opened up a wide range of real-world applications, including movie production [65], gaming [48], and even serving as virtual environments for robotics and embodied AI learning [2, 3]. Despite the great visual quality, accurately simulating the correct physical-related effects remains challenging and is often overlooked by existing methods [25, 61]. These models learn physics by exhaustively training the model over a large corpus of video data. The complexity of real-world motion, interactions, and dynamics make such implicit learning difficult and less effective [7].
@@
In contrast, traditional graphics approaches [19, 20, 21, 39, 34] predefine physical parameters for a scene using simulation engines, enabling them to render results that are inherently physically accurate. However, manually specifying these physical properties is often labor-intensive and becomes impractical as scene complexity increases. As a result, such methods are typically constrained to simple physical scenarios, such as perfectly elastic collisions [25] and basic rigid body dynamics [34]. The capbility to scale to complex real-world environments, where physical effects emerge from intricate motion, interactions, forces, collisions, and environmental context, remains an open challenge.
@@
In this work, we propose DiffPhy, a novel framework that enables video diffusion models to generate physically accurate and visually compelling videos from arbitrary user prompts. At the heart of our approach is the use of Large Language Models (LLMs) [1, 7] as both reasoners and evaluators to guide the generation process. Specifically, our method begins by using an LLM [1] to infer a rich physical context from the input text. This inferred context is then integrated into the prompt, yielding an enhanced text instruction that provides explicit physical cues for video generation. We then fine-tune a pre-trained video diffusion model [50] using this enhanced prompt to make it physics-aware. During fine-tuning, a Multimodal Large Language Model (MLLM) [7] evaluates the generated video by analyzing their physical alignment and semantic consistency with the text description as well as their physical plausibility. This supervision is implemented as a set of novel training objectives that enforce both physical correctness and semantic consistency with the original prompt.
@@
One challenge left is how to obtain a high-quality, annotated video dataset that captures diverse physical scenarios for effective fine-tuning. Most existing datasets are designed primarily for evaluation [6, 7, 34, 38], and typically consist of synthetic or simulated scenes with limited scale, diversity, and complexity. Fine-tuning directly on such datasets often leads to degraded performance. To this end, we also introduce a larger curated dataset of real videos that covers diverse entities, motions, behaviors and physical phenomena. This dataset can be effectively used for training video generation model, boosting their capabilities to handle complex real-world physical scenarios.
@@ 2 Related Work
Text-to-Video Generation. Diffusion-based text-to-video generation is undergoing a rapid revolution, with improvements in quality [2, 63, 8], controllability [67, 4, 44], efficiency [58, 66, 24], extrapolation [68, 11, 36], etc. Incorporating pre-trained text-to-image diffusion models, early works, such as AnimateDiff [15] and SVD [8], modify the UNet blocks to model the temporal dependencies. Recent advancements [63, 27, 51] leverage the scalable Diffusion Transformer [41] (DiT) with more parameters and training data and present better synthesis quality and diversity. Closed-source models, such as SORA [40], Pika [42], and Kling [26], have demonstrated their impressive capabilities in video production. However, these existing methods implicitly learn the physical principles from data distribution and often neglect real-world physical laws [7], focusing more on visual aspects like color and texture rather than underlying physical dynamics [25]. Consequently, the generated videos frequently suffer from a lack of physical commonsense and fail to capture authentic object motion.
Video Physics Reasoning. Existing work has used physics simulators and engines for tasks such as rigid body dynamics, fluid simulation, and deformable object modeling [19, 20, 21, 39, 34]. Recent methods [12, 29, 30, 56, 55, 59, 60] extract physical representations from visual inputs via neural networks for downstream reasoning or simulation. While effective in domains like graphics, robotics, and scientific computing, these approaches often rely on rule-based, handcrafted solvers, which limits their expressiveness, scalability, and requiring extensive manual tuning. Moreover, they are typically confined to synthetic data and can not scale to complex, real-world video generation.
LLM in T2V Generation. LLMs have been widely used for prompt refinement in text-to-image and text-to-video generation, helping interpret prompts and guide layout planning [9, 22, 31, 33, 57, 62, 69]. Methods like PhyT2V [61] refine prompts iteratively, but they operate solely in the text domain and do not improve the generation model itself. Recent work has proposed physical evaluation benchmarks that assess physical plausibility using multimodal LLMs [7, 6, 37], leveraging models such as VideoCon [5], GPT-4o [1], LLaVA [32], and InternVideo [53]. We adopt the multimodal LLM evaluator VideoCon-Physics [7] for its lightweight, open-source design. It is built on VideoCon [5], with CLIP [43] as the visual encoder and LLaMA-7B [47] as the language backbone.
@@ 3 Method
We present DiffPhy, a generic framework that enables diffusion models to generate physically-correct and photo-realistic videos. Our core idea is to (1) ground the user prompt with rich physical context through LLM reasoning and (2) fine-tuning the diffusion model to make it physics-aware by ensuring the generated physical effects are valid. An overview of our method is shown in Figure 2. In more detail, given a user prompt, we first apply a pretrained LLM (e.g. GPT-4) to reason the underlying physical attributes (e.g. gravity, velocity, collisions, bounce) via Chain-of-Thought [54] prompting (Figure 2-(1)). Then, based on the inferred properties, we query the LLM again to enhance text prompt to include physical context (Figure 2-(2a)). The enhanced text description is then fed to the video diffusion model to guide the generation process (Figure 2-(3)). We also request the LLM to generate a list of physical phenomena (Figure 2-(2b)) associated with the described event. During training, this list is used to penalize any synthesized videos that contain invalid physical effects (Figure 2-(4)). We make the diffusion model physics-aware by training with three novel loss terms that jointly encourage the physical correctness and video-realism. In the following, we first describe how to obtain enhanced text prompt with detailed physical attributes (Section 3.1) and then discuss how to train the diffusion model to produce physically-correct videos (Section 3.2).
@@ 3.1 Grounding Text Description with Physical Context
As shown in Figure 2, user prompts are often simple and incomplete, without detailed descriptions about the physical effects associated with the event/action. Existing methods [50, 40, 49, 63, 8] implicitly interpret physics from such text by exhaustively training over a large data corpus, which is known to be difficult and costly [7]. Instead, our method seeks to explicitly mine the physical attributes/effects in the user prompt as a pre-processing step. Thereafter, the generation process can be guided by an enhanced text prompt that contains sufficient physical context. Specifically, we leverage a pre-trained LLM to reason physical context and obtain three types of text descriptions: (1) a list of physical attributes that capture roles (cause/effect), interactions, and physics; (2) an enhanced prompt with scene and physical details; (3) a list of physical phenomena associated with the event.
@@
To extract physical attributes from text, we prompt a pretrained LLM using chain-of-thought (CoT) [54]. Specifically, we prompt the language model by “list all relevant physical principles, explain their roles, and formalize them as quantitative constraints." This allow the LLM to effectively identify event semantics including underlying forces (e.g., gravity, friction), kinematic relationships (e.g., constant velocity, acceleration), and interaction rules (e.g., elastic vs. inelastic collisions). We then parse the LLM output into structured machine-readable rules, creating a coherent physical description of the target scene. In practice, the user prompt might not be complete and can sometime omit important entities or interactions. For example, an incomplete text prompt might like “A balloon is popped by being over-inflated, causing it to burst", which does not specify the cause for over-inflation. As a result, a balloon that pops without cause becomes an invalid scene. To address this, we extend the CoT reasoning into a procedural process through four detailed steps: (1) identifying the key physical principles involved; (2) determining which entities initiate actions and which are affected; (3) reasoning about how these entities interact; and (4) noting any resulting physical phenomena. As a result, we obtain a list of physical attributes that are both comprehensive and valid. The detailed prompt is provided in Appendix.
@@ 3.2 Physics-aware Model Training
One challenge left is how to train a diffusion model to effectively leverage the rich physical context in the text description as well as precisely generate the desired effects. This indicates that the diffusion model needs to (1) comply with specific physical rules implied in the text prompt; (2) respect the commonsense as a natural video and (3) produce desired content adhering to the text. Ideally, this can be accomplished by optimizing the diffusion model towards some metrics that evaluate how well a generated video is aligned with specific physical effects. Yet to the best of our knowledge, there is no such measurement in the literature. To bridge this gap, we propose to adopt a multi-modal LLM (MLLM) [7] as a learned evaluator, learning its strong capability to perform complex physical reasoning over video and text. Specifically, We design three MLLM-based losses, i.e., physical phenomena loss, commonsense loss, semantic consistency loss, alongside the standard denoising score matching objective [16]. These losses are computed by querying MLLM over generated videos.
@@
Specifically, during training, at a sampled timestep $t$, we decode the predicted latent clip $x_{t}\in R^{m\times c\times h\times w}$ with $m$ frames into a pixel space video $v_{t}$. Then, we employ an MLLM [7] to evaluate the alignment between $v_{t}$ and each fact $f_{i}$ in the physical phenomena $\mathbf{F}=\{f_{1},...,f_{n}\}$, one by one. For each fact $f_{i}$, the MLLM outputs an integer label $l_{i}\in\{0,1\}$, indicating the fact is ‘"not matched" or "matched" in $v_{t}$. This allows us to define a loss function to encourage the diffusion model to produce "matched" results and penalize "unmatched" results. Specifically, for a fact $f_{i}$, we can compute an expected label $\mathbb{E}(l_{i})$ by weighted averaging over all labels. Formally,
@@
where $z_{k}$ corresponds the logit associated with the word $k$ predicted by the MLLM. Given the expected label, we then define the loss function as $\mathcal{L_{\text{phen}}}=\sum_{f_{i}\in\mathbf{F}}||\mathbb{E}(l_{i})-1||_{2}^{2}$. Training the model with $\mathcal{L_{\text{phen}}}$ encourages the diffusion model to generate the desired physical effects.
@@
It is equally important for the diffusion model to resolve any failed or incorrect facts. For a "not matched" fact $f_{i}$, we introduce an additional module to guide the diffusion model to better understand the failure or missed cases during generation. This module encodes $f_{i}$ into text embeddings and conditions the diffusion model via a separate cross-attention layer in each DIT block, as shown in Figure 2-(3). To stablize the training, we copy the original cross-attention module and connect it with the original model via a zero-initialized convolution layer (ZeroConv) in a way similar to [64]. To reduce training cost, we apply a LoRA adapter [18] to the cloned attention module and fine-tune only the adapter. More implementation details can be found in the Appendix.
@@
We also introduce a physical commonsense loss $\mathcal{L}_{\text{com}}$ to evaluate the overall physical plausibility of the scene. We use the same multi-modal LLM [7] again to assign an integer-valued plausibility score $s_{com}$ to $v_{t}$, ranging from 1 to 5, where 5 indicates the highest plausibility. Similar to $\mathcal{L}_{phen}$, we compute an expected score $s_{\text{com}}$ over all possible values (i.e. {1,2,3,4,5}) using Eq. 1. Then, the physical commonsense loss can be defined as a mean squared error (MSE): $\mathcal{L}_{com}=\|\mathbb{E}(s_{\text{com}})/\tau-1\|_{2}^{2}$, where $\tau=5$ is a normalization constant. This loss encourages the diffusion model to generate videos that meet physical commonsense.
@@
To ensure the generated content still follows the input prompt, we introduce a semantic consistency loss. Again, we use the same pretrained multi-modal LLM [7] to give scores from 1 to 5 for evaluating the semantic consistency between $v_{t}$ and the input text prompt. The semantic consistency loss has the same format as physical commonsense loss, i.e. $\mathcal{L}_{sem}=\|\mathbb{E}(s_{\text{sem}})/\tau-1\|_{2}^{2}$, where $\tau=5$ is a normalization constant, $s_{\text{sem}}$ is the expected score over all possible values (i.e. {1,2,3,4,5}).
@@
The final loss is computed by combining the original denoising score matching $\mathcal{L}_{\epsilon}$ objective [16] with the three newly proposed MLLM-based loss functions, i.e., phenomena loss $\mathcal{L}_{\text{phen}}$, physical commensense loss $\mathcal{L}_{\text{com}}$ and semantic consistency loss $\mathcal{L}_{\text{sem}}$. Formally,

### v2

@@ Abstract
Figure 1: DiffPhy enables physically grounded and semantically aligned video generation across diverse real-world scenarios, including gravity-driven motion, fluid interactions, forceful impacts, and object manipulation. It outperforms state-of-the-art video diffusion model Wan 2.1–14B wan14b in both visual plausibility and physical coherence.
@@ 1 Introduction
Recent advances in video diffusion models have made it possible to generate high-fidelity videos with rich details from text prompts guo2024animatediff ; wan14b ; yang2024cogvideox ; chen2024videocrafter2 ; kong2024hunyuanvideo ; agarwal2025cosmos ; openai2024sora ; lumalabsLumaRay2 ; wang2023laviehighqualityvideogeneration ; fan2025vchitect ; pika2025 ; gen32024 ; klingai . Such generation capability has opened up a wide range of real-world applications, including movie production zhang2025generative , gaming valevski2024diffusion , and even serving as virtual environments for robotics and embodied AI learning agarwal2025cosmos ; alhaija2025cosmos . Despite the great visual quality, accurately simulating the correct physical-related effects remains challenging and is often overlooked by existing methods kang2024far ; xue2024phyt2v . These models learn physics by exhaustively training the model over a large corpus of video data. The complexity of real-world motion, interactions, and dynamics make such implicit learning difficult and less effective bansal2025videophy .
@@
In contrast, traditional graphics approaches hu2019difftaichi ; hu2018moving ; hu2019taichi ; nvidia2019physx ; liu2024physgen predefine physical parameters for a scene using simulation engines, enabling them to render results that are inherently physically accurate. However, manually specifying these physical properties is often labor-intensive and becomes impractical as scene complexity increases. As a result, such methods are typically constrained to simple physical scenarios, such as perfectly elastic collisions kang2024far and basic rigid body dynamics liu2024physgen . The capbility to scale to complex real-world environments, where physical effects emerge from intricate motion, interactions, forces, collisions, and environmental context, remains an open challenge.
@@
In this work, we propose DiffPhy, a novel framework that enables video diffusion models to generate physically accurate and visually compelling videos from arbitrary user prompts. At the heart of our approach is the use of Large Language Models (LLMs) achiam2023gpt4 ; bansal2025videophy as both reasoners and evaluators to guide the generation process. Specifically, our method begins by using an LLM achiam2023gpt4 to infer a rich physical context from the input text. This inferred context is then integrated into the prompt, yielding an enhanced text instruction that provides explicit physical cues for video generation. We then fine-tune a pre-trained video diffusion model wan14b using this enhanced prompt to make it physics-aware. During fine-tuning, a Multimodal Large Language Model (MLLM) bansal2025videophy evaluates the generated video by analyzing their physical alignment and semantic consistency with the text description as well as their physical plausibility. This supervision is implemented as a set of novel training objectives that enforce both physical correctness and semantic consistency with the original prompt.
@@
One challenge left is how to obtain a high-quality, annotated video dataset that captures diverse physical scenarios for effective fine-tuning. Most existing datasets are designed primarily for evaluation bansal2024videophy ; bansal2025videophy ; liu2024physgen ; motamed2025generative , and typically consist of synthetic or simulated scenes with limited scale, diversity, and complexity. Fine-tuning directly on such datasets often leads to degraded performance. To this end, we also introduce a larger curated dataset of real videos that covers diverse entities, motions, behaviors and physical phenomena. This dataset can be effectively used for training video generation model, boosting their capabilities to handle complex real-world physical scenarios.
@@ 2 Related Work
Text-to-Video Generation. Diffusion-based text-to-video generation is undergoing a rapid revolution, with improvements in quality agarwal2025cosmos ; yang2024cogvideox ; blattmann2023stable , controllability zhang2025enabling ; bai2025recammaster ; ren2025gen3c , efficiency xi2025sparse ; zhang2025flashvideo ; kahatapitiya2024adaptive , extrapolation zhao2025riflex ; dalal2025one ; mao2024story , etc. Incorporating pre-trained text-to-image diffusion models, early works, such as AnimateDiff guo2024animatediff and SVD blattmann2023stable , modify the UNet blocks to model the temporal dependencies. Recent advancements yang2024cogvideox ; kong2024hunyuanvideo ; wang2025wan leverage the scalable Diffusion Transformer peebles2023scalable (DiT) with more parameters and training data and present better synthesis quality and diversity. Closed-source models, such as SORA openai2024sora , Pika pika2025 , and Kling klingai , have demonstrated their impressive capabilities in video production. However, these existing methods implicitly learn the physical principles from data distribution and often neglect real-world physical laws bansal2025videophy , focusing more on visual aspects like color and texture rather than underlying physical dynamics kang2024far . Consequently, the generated videos frequently suffer from a lack of physical commonsense and fail to capture authentic object motion.
Video Physics Reasoning. Existing work has used physics simulators and engines for tasks such as rigid body dynamics, fluid simulation, and deformable object modeling hu2019difftaichi ; hu2018moving ; hu2019taichi ; nvidia2019physx ; liu2024physgen . Recent methods davis2015visual ; li2023pac ; li2020visual ; wu2017learning ; wu2016physics ; xia2024video2game ; xu2019densephysnet extract physical representations from visual inputs via neural networks for downstream reasoning or simulation. While effective in domains like graphics, robotics, and scientific computing, these approaches often rely on rule-based, handcrafted solvers, which limits their expressiveness, scalability, and requiring extensive manual tuning. Moreover, they are typically confined to synthetic data and can not scale to complex, real-world video generation.
LLM in T2V Generation. LLMs have been widely used for prompt refinement in text-to-image and text-to-video generation, helping interpret prompts and guide layout planning carreira2018short ; huang2025modality ; lian2023llm ; lin2023videodirectorgpt ; wu2024self ; yang2024mastering ; zhu2024compositional . Methods like PhyT2V xue2024phyt2v refine prompts iteratively, but they operate solely in the text domain and do not improve the generation model itself. Recent work has proposed physical evaluation benchmarks that assess physical plausibility using multimodal LLMs bansal2025videophy ; bansal2024videophy ; meng2024towards , leveraging models such as VideoCon bansal2024videocon , GPT-4o achiam2023gpt4 , LLaVA lin2023videoLLaVA , and InternVideo wang2024internvideo2 . We adopt the multimodal LLM evaluator VideoCon-Physics bansal2025videophy for its lightweight, open-source design. It is built on VideoCon bansal2024videocon , with CLIP radford2021learning as the visual encoder and LLaMA-7B touvron2023llama as the language backbone.
@@ 3 Method
We present DiffPhy, a generic framework that enables diffusion models to generate physically-correct and photo-realistic videos. Our core idea is to (1) ground the user prompt with rich physical context through LLM reasoning and (2) fine-tuning the diffusion model to make it physics-aware by ensuring the generated physical effects are valid. An overview of our method is shown in Figure 2. In more detail, given a user prompt, we first apply a pretrained LLM (e.g. GPT-4) to reason the underlying physical attributes (e.g. gravity, velocity, collisions, bounce) via Chain-of-Thought wei2022chain prompting (Figure 2-(1)). Then, based on the inferred properties, we query the LLM again to enhance text prompt to include physical context (Figure 2-(2a)). The enhanced text description is then fed to the video diffusion model to guide the generation process (Figure 2-(3)). We also request the LLM to generate a list of physical phenomena (Figure 2-(2b)) associated with the described event. During training, this list is used to penalize any synthesized videos that contain invalid physical effects (Figure 2-(4)). We make the diffusion model physics-aware by training with three novel loss terms that jointly encourage the physical correctness and video-realism. In the following, we first describe how to obtain enhanced text prompt with detailed physical attributes (Section 3.1) and then discuss how to train the diffusion model to produce physically-correct videos (Section 3.2).
@@ 3.1 Grounding Text Description with Physical Context
As shown in Figure 2, user prompts are often simple and incomplete, without detailed descriptions about the physical effects associated with the event/action. Existing methods wan14b ; openai2024sora ; deepmindveo2 ; yang2024cogvideox ; blattmann2023stable implicitly interpret physics from such text by exhaustively training over a large data corpus, which is known to be difficult and costly bansal2025videophy . Instead, our method seeks to explicitly mine the physical attributes/effects in the user prompt as a pre-processing step. Thereafter, the generation process can be guided by an enhanced text prompt that contains sufficient physical context. Specifically, we leverage a pre-trained LLM to reason physical context and obtain three types of text descriptions: (1) a list of physical attributes that capture roles (cause/effect), interactions, and physics; (2) an enhanced prompt with scene and physical details; (3) a list of physical phenomena associated with the event.
@@
To extract physical attributes from text, we prompt a pretrained LLM using chain-of-thought (CoT) wei2022chain . Specifically, we prompt the language model by “list all relevant physical principles, explain their roles, and formalize them as quantitative constraints." This allow the LLM to effectively identify event semantics including underlying forces (e.g., gravity, friction), kinematic relationships (e.g., constant velocity, acceleration), and interaction rules (e.g., elastic vs. inelastic collisions). We then parse the LLM output into structured machine-readable rules, creating a coherent physical description of the target scene. In practice, the user prompt might not be complete and can sometime omit important entities or interactions. For example, an incomplete text prompt might like “A balloon is popped by being over-inflated, causing it to burst", which does not specify the cause for over-inflation. As a result, a balloon that pops without cause becomes an invalid scene. To address this, we extend the CoT reasoning into a procedural process through four detailed steps: (1) identifying the key physical principles involved; (2) determining which entities initiate actions and which are affected; (3) reasoning about how these entities interact; and (4) noting any resulting physical phenomena. As a result, we obtain a list of physical attributes that are both comprehensive and valid. The detailed prompt is provided in Appendix.
@@ 3.2 Physics-aware Model Training
One challenge left is how to train a diffusion model to effectively leverage the rich physical context in the text description as well as precisely generate the desired effects. This indicates that the diffusion model needs to (1) comply with specific physical rules implied in the text prompt; (2) respect the commonsense as a natural video and (3) produce desired content adhering to the text. Ideally, this can be accomplished by optimizing the diffusion model towards some metrics that evaluate how well a generated video is aligned with specific physical effects. Yet to the best of our knowledge, there is no such measurement in the literature. To bridge this gap, we propose to adopt a multi-modal LLM (MLLM) bansal2025videophy as a learned evaluator, learning its strong capability to perform complex physical reasoning over video and text. Specifically, We design three MLLM-based losses, i.e., physical phenomena loss, commonsense loss, semantic consistency loss, alongside the standard denoising score matching objective ho2020denoising . These losses are computed by querying MLLM over generated videos.
@@
Specifically, during training, at a sampled timestep $t$, we decode the predicted latent clip $x_{t}\in R^{m\times c\times h\times w}$ with $m$ frames into a pixel space video $v_{t}$. Then, we employ an MLLM bansal2025videophy to evaluate the alignment between $v_{t}$ and each fact $f_{i}$ in the physical phenomena $\mathbf{F}=\{f_{1},...,f_{n}\}$, one by one. For each fact $f_{i}$, the MLLM outputs an integer label $l_{i}\in\{0,1\}$, indicating the fact is ‘"not matched" or "matched" in $v_{t}$. This allows us to define a loss function to encourage the diffusion model to produce "matched" results and penalize "unmatched" results. Specifically, for a fact $f_{i}$, we can compute an expected label $\mathbb{E}(l_{i})$ by weighted averaging over all labels. Formally,
@@
[…]
@@
It is equally important for the diffusion model to resolve any failed or incorrect facts. For a "not matched" fact $f_{i}$, we introduce an additional module to guide the diffusion model to better understand the failure or missed cases during generation. This module encodes $f_{i}$ into text embeddings and conditions the diffusion model via a separate cross-attention layer in each DIT block, as shown in Figure 2-(3). To stablize the training, we copy the original cross-attention module and connect it with the original model via a zero-initialized convolution layer (ZeroConv) in a way similar to zhang2023adding . To reduce training cost, we apply a LoRA adapter hu2022lora to the cloned attention module and fine-tune only the adapter. More implementation details can be found in the Appendix.
@@
We also introduce a physical commonsense loss $\mathcal{L}_{\text{com}}$ to evaluate the overall physical plausibility of the scene. We use the same multi-modal LLM bansal2025videophy again to assign an integer-valued plausibility score $s_{com}$ to $v_{t}$, ranging from 1 to 5, where 5 indicates the highest plausibility. Similar to $\mathcal{L}_{phen}$, we compute an expected score $s_{\text{com}}$ over all possible values (i.e. {1,2,3,4,5}) using Eq. 1. Then, the physical commonsense loss can be defined as a mean squared error (MSE): $\mathcal{L}_{com}=\|\mathbb{E}(s_{\text{com}})/\tau-1\|_{2}^{2}$, where $\tau=5$ is a normalization constant. This loss encourages the diffusion model to generate videos that meet physical commonsense.
@@
To ensure the generated content still follows the input prompt, we introduce a semantic consistency loss. Again, we use the same pretrained multi-modal LLM bansal2025videophy to give scores from 1 to 5 for evaluating the semantic consistency between $v_{t}$ and the input text prompt. The semantic consistency loss has the same format as physical commonsense loss, i.e. $\mathcal{L}_{sem}=\|\mathbb{E}(s_{\text{sem}})/\tau-1\|_{2}^{2}$, where $\tau=5$ is a normalization constant, $s_{\text{sem}}$ is the expected score over all possible values (i.e. {1,2,3,4,5}).
@@
The final loss is computed by combining the original denoising score matching $\mathcal{L}_{\epsilon}$ objective ho2020denoising with the three newly proposed MLLM-based loss functions, i.e., phenomena loss $\mathcal{L}_{\text{phen}}$, physical commensense loss $\mathcal{L}_{\text{com}}$ and semantic consistency loss $\mathcal{L}_{\text{sem}}$. Formally,